OpenStudy (anonymous):

22x - y = -23 ( equation 1) 11x + 12y = 1 ( equation 2) To find the value of x , you need to substitute _____ in equation 2. The value of x is ?

OpenStudy (anonymous):

you can solve the first one for \(y\) and get \[y=22x+23\] substitute that

OpenStudy (anonymous):

X = ( y-23)/22 X= ( y + 23)/22 Y= 22x + 23 Y= 22x - 23

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's the answer choices

OpenStudy (anonymous):

? Anyone

OpenStudy (wolf1728):

The easiest way to solve it would be to multiply equation 2 by -2 11x + 12y = 1 * -2 = -22x -24y -=2 Now add this to equation1 22x - y = -23 ( equation 1) -22x -24y -=2 ( equation 2) Adding the equations we get: -25y = -25 y=1 Substituting this into equation 1 we get: 22x -1 = -23 22x = -22 x = -1